OpenStudy (anonymous):

Choose the fraction in simplest form that correctly expresses this ratio. 84:112

OpenStudy (alexalas):

may we see the list also do you know what a least common denominator is?

OpenStudy (anonymous):

3/4 1/3 2/3

OpenStudy (alexalas):

there are two ways to do this First way 84/112= A 3/4=.75 B 1/3=.3333333 C 2/3=.6666666